# Curator: Efficient Vector Search with Low-Selectivity Filters

### v2

@@ Abstract.
copyright: none
@@ 1. Introduction
The rise of deep learning has transformed information retrieval by enabling the representation of unstructured data as high-dimensional vectors that capture semantic relationships (Le and Mikolov, 2014; Radford et al., 2021; Narayanan et al., 2017). Approximate nearest-neighbor search (ANNS) over these vector embeddings has become a fundamental primitive powering modern applications from web search (Nayak, 2019; Zeng, 2022) to recommendation systems (Zhang et al., 2019; Naumov et al., 2019; Covington et al., 2016) and retrieval-augmented generation (OpenAI, 2023; Lewis et al., 2020).
@@ 4.1. Single-Label Search
Input : Index $T$, query vector $x_{q}$, label $l$, beam width $b$, result set size efef
Output : $k$ approximate nearest neighbors of $x_{q}$
4
Initialize $\mathcal{Q}\leftarrow\textnormal{{BeamSearch}}(T.root,b)$
@@
5
6while $\mathcal{Q}\neq\varnothing$ do
@@ 13    $\mathcal{R}^{\prime}\leftarrow\mathcal{R}\cup\mathcal{P}_{n,l}$
14    if $|\mathcal{R}^{\prime}|>ef$ then
15       $\mathcal{R}^{\prime}\leftarrow$ top-ef closest vectors in $\mathcal{R}^{\prime}$
16
@@ 23    $\mathcal{Q}\leftarrow\mathcal{Q}\cup C_{n}$
Phase 2: Best-First Search. Using the frontier initialized by beam search, the second phase performs best-first search, systematically exploring the neighborhood guided by heuristic scores. The algorithm processes nodes based on their type with regard to the per-label index $T_{l}$: leaf nodes containing buffers update the result set with their qualified vectors. For nodes without buffers, the Bloom filter $\mathcal{B}_{n}$ determines whether the node is an internal node requiring expansion or an external node outside $T_{l}$ where the search should backtrack. During exploration, the search maintains a dynamic result set $\mathcal{R}$ containing the $ef$ closest vectors discovered so far. Each time a buffer is visited, the algorithm updates $\mathcal{R}$ and checks for convergence—if the result set remains unchanged (i.e., no vector in the buffer is closer than the worst member of $\mathcal{R}$), the search terminates. Larger $ef$ values widen the exploration radius, trading more work for higher recall. In the limit, when $ef$ is sufficiently large, the traversal visits the entire per-label index encompassing all qualifying vectors, thus guaranteeing completeness. We found this adaptive termination criterion accommodates varying query difficulties more effectively than approaches that use a fixed computational budget.
@@ 11 $n_{\sigma}\leftarrow$ InitializeNode()
    /* $c$ covers vector ID range $[c.min,c.max)$ */
@@
Temporary Index Construction. Algorithm 2 presents the temporary index construction procedure. The algorithm exploits the sorted vector ID property through a recursive tree building process that mirrors the structure of the main index $T$. The BuildSubtree function constructs each node $n_{\sigma}$ in the temporary index $T_{\sigma}$ by processing a corresponding range $\mathcal{P}_{\sigma}[l:r]$ of sorted qualified vector IDs. The algorithm operates as follows for each node: (1) Leaf condition: If the number of qualified vectors in the current range does not exceed the buffer capacity $B_{max}$, the algorithm creates a leaf node and terminates recursion for this sub-tree. (2) Hierarchical partitioning: Otherwise, it iterates through each child $c$ of the corresponding main index node $n$. For each child, binary search operations locate the subset of qualified vectors whose IDs fall within the child’s vector ID range $[c.min,c.max)$. This partitioning step leverages the contiguous property of sorted vector IDs to efficiently split the qualified vector list. After partitioning, the construction algorithm is then applied recursively to build the sub-tree for each child.
Figure 5 illustrates this process using a simplified scenario where each node has two children and both leaf capacity and buffer capacity are 2, allowing vector IDs to be represented as 3-digit binary numbers. The left panel shows the base index with vector ID ranges annotated for each node, such as $000\text{-}111]$ for the root and $000\text{-}011]$, $100\text{-}111]$ for its children, corresponding to the $[c.min,c.max)$ ranges referenced in the algorithm. The middle panel demonstrates the recursive partitioning process: starting with sorted qualified vector IDs $\{000,010,100,101,110\}$, binary search finds the partition boundary at vector ID $100$ (the start of the right sub-tree’s range $100\text{-}111]$), splitting the list into groups $\{000,010\}$ and $\{100,101,110\}$. The partitioning continues recursively until the leaf condition is met—for instance, the right sub-tree is further split into $\{100,101\}$ and $\{110\}$, with the latter qualifying the leaf condition ($\leq 2$ vectors). The right panel shows the resulting temporary index where the qualified vector IDs are distributed into buffers in leaf nodes.
@@ 6.1. Experimental Setup
In addition to these standalone vector indexes, we evaluate pgvector (pgvector contributors, 2024), a PostgreSQL extension that allows users to create vector-typed columns and build vector indexes on them, integrating vector search with SQL-based predicate filtering. We evaluate two pgvector-based baselines: Pg-IVF and Pg-HNSW, which use IVF-Flat and HNSW as the underlying vector indexes, respectively. Vectors and labels are stored in PostgreSQL tables with labels represented as INT[] arrays indexed using GIN for efficient containment queries. We use UNLOGGED tables to reduce index construction time. For Pg-IVF, we scan all vectors in the nprobenprobe buckets closest to the query vector and return the top-$k$ qualified vectors. Distances are computed for both qualified and unqualified vectors. Pg-HNSW employs iterative search that resumes exploration from the next $ef$ best candidates after each search iteration converges, terminating after finding $k$ qualified vectors or scanning at most $T_{max}$ vectors. To generate recall-throughput curves, we sweep nprobenprobe for Pg-IVF and $T_{max}$ for Pg-HNSW.
@@ 6.7. Ablation Study
Structural Constraints. Curator enforces structural sharing by constraining per-label indexes to follow the base index’s hierarchical partitioning, which potentially impacts performance since per-label indexes cannot freely adapt to label distributions. To evaluate this trade-off, we compare Curator against a variant where per-label indexes are constructed independently with the same hyper-parameters. Fig. 15a shows nearly identical performance between both versions, demonstrating that structural constraints have minimal impact on the quality of embedded per-label indexes. This finding is further supported by the similar performance between Curator with predicate indexed versus per-predicate indexing in Figure 9.
Vector-Label Correlation. Real-world datasets exhibit varying degrees of vector-label correlation—the tendency for vectors sharing the same label to cluster together in the embedding space due to semantic similarity. To assess Curator’s robustness across datasets with different correlation patterns, we test on a shuffled variant of YFCC-10M where labels are randomly reassigned, effectively eliminating vector-label correlation. Fig. 15b shows nearly identical recall-throughput trade-offs on both original and shuffled datasets, confirming Curator’s robust performance across varied data distributions. We additionally quantify correlation using the ”Query Correlation” metric from ACORN (Patel et al., 2024), which compares a query’s distance to the set of qualified vectors against a random set of equal size; larger values indicate stronger clustering by label. On YFCC-10M, the average query correlation is 0.161, whereas on the label-shuffled dataset it is 0.028, showing the effectiveness of label shuffling and confirming that the original dataset exhibits meaningful correlation.
Tree Traversal Overhead. Unlike graph-based indexes where all distance computations target individual vectors, Curator performs distance computations against both cluster centroids (during tree traversal) and vectors (during buffer scanning). To quantify the overhead of tree traversal, we measure the proportion of distance computations used for tree traversal versus buffer scanning. Fig. 15c shows that traversal overhead decreases from 26.9% to 18.9% as dataset size grows from 2M to 10M vectors, demonstrating that the overhead remains moderate and diminishes at larger scales.

### v3

@@ Abstract.
copyright: none11footnotetext: Yongji Wu and Xiao Zhang were with Duke University at the time this work was conducted.
@@ 1. Introduction
The rise of deep learning has transformed information retrieval by enabling the representation of unstructured data as high-dimensional vectors that capture semantic relationships (Le and Mikolov, 2014; Radford et al., 2021; Narayanan et al., 2017). Approximate nearest neighbor search (ANNS) over these vector embeddings has become a fundamental primitive powering modern applications from web search (Nayak, 2019; Zeng, 2022) to recommendation systems (Zhang et al., 2019; Naumov et al., 2019; Covington et al., 2016) and retrieval-augmented generation (OpenAI, 2023; Lewis et al., 2020).
@@ 4.1. Single-Label Search
Input : Index $T$, query vector $x_{q}$, label $l$, beam width $b$, result set size $\mathit{ef}$
Output : $k$ approximate nearest neighbors of $x_{q}$
@@
Initialize $\mathcal{Q}\leftarrow\textnormal{{BeamSearch}}(T.\mathit{root},b)$
@@
4
5while $\mathcal{Q}\neq\varnothing$ do
@@ 12    $\mathcal{R}^{\prime}\leftarrow\mathcal{R}\cup\mathcal{P}_{n,l}$
13    if $|\mathcal{R}^{\prime}|>\mathit{ef}$ then
14       $\mathcal{R}^{\prime}\leftarrow$ top-$\mathit{ef}$ closest vectors in $\mathcal{R}^{\prime}$
15
@@ 22    $\mathcal{Q}\leftarrow\mathcal{Q}\cup C_{n}$
Phase 2: Best-First Search. Using the frontier initialized by beam search, the second phase performs best-first search, systematically exploring the neighborhood guided by heuristic scores. The algorithm processes nodes based on their type with regard to the per-label index $T_{l}$: leaf nodes containing buffers update the result set with their qualified vectors. For nodes without buffers, the Bloom filter $\mathcal{B}_{n}$ determines whether the node is an internal node requiring expansion or an external node outside $T_{l}$ where the search should backtrack. During exploration, the search maintains a dynamic result set $\mathcal{R}$ containing the $\mathit{ef}$ closest vectors discovered so far. Each time a buffer is visited, the algorithm updates $\mathcal{R}$ and checks for convergence—if the result set remains unchanged (i.e., no vector in the buffer is closer than the worst member of $\mathcal{R}$), the search terminates. Larger $\mathit{ef}$ values widen the exploration radius, trading more work for higher recall. In the limit, when $\mathit{ef}$ is sufficiently large, the traversal visits the entire per-label index encompassing all qualifying vectors, thus guaranteeing completeness. We found this adaptive termination criterion accommodates varying query difficulties more effectively than approaches that use a fixed computational budget.
@@ 11 $n_{\sigma}\leftarrow$ InitializeNode()
    /* $c$ covers vector ID range $c.\mathit{min},c.\mathit{max}) *$/
@@
Temporary Index Construction. Algorithm 2 presents the temporary index construction procedure. The algorithm exploits the sorted vector ID property through a recursive tree building process that mirrors the structure of the main index $T$. The BuildSubtree function constructs each node $n_{\sigma}$ in the temporary index $T_{\sigma}$ by processing a corresponding range $\mathcal{P}_{\sigma}[l:r]$ of sorted qualified vector IDs. The algorithm operates as follows for each node: (1) Leaf condition: If the number of qualified vectors in the current range does not exceed the buffer capacity $B_{max}$, the algorithm creates a leaf node and terminates recursion for this sub-tree. (2) Hierarchical partitioning: Otherwise, it iterates through each child $c$ of the corresponding main index node $n$. For each child, binary search operations locate the subset of qualified vectors whose IDs fall within the child’s vector ID range $c.\mathit{min},c.\mathit{max})$. This partitioning step leverages the contiguous property of sorted vector IDs to efficiently split the qualified vector list. After partitioning, the construction algorithm is then applied recursively to build the sub-tree for each child.
Fig. 5 illustrates this process using a simplified scenario where each node has two children and both leaf capacity and buffer capacity are 2, allowing vector IDs to be represented as 3-digit binary numbers. The left panel shows the base index with vector ID ranges annotated for each node, such as $000\text{-}111]$ for the root and $000\text{-}011]$, $100\text{-}111]$ for its children, corresponding to the $c.\mathit{min},c.\mathit{max})$ ranges referenced in the algorithm. The middle panel demonstrates the recursive partitioning process: starting with sorted qualified vector IDs $\{000,010,100,101,110\}$, binary search finds the partition boundary at vector ID $100$ (the start of the right sub-tree’s range $100\text{-}111]$), splitting the list into groups $\{000,010\}$ and $\{100,101,110\}$. The partitioning continues recursively until the leaf condition is met—for instance, the right sub-tree is further split into $\{100,101\}$ and $\{110\}$, with the latter qualifying the leaf condition ($\leq 2$ vectors). The right panel shows the resulting temporary index where the qualified vector IDs are distributed into buffers in leaf nodes.
@@ 6.1. Experimental Setup
In addition to these standalone vector indexes, we evaluate pgvector (pgvector contributors, 2024), a PostgreSQL extension that allows users to create vector-typed columns and build vector indexes on them, integrating vector search with SQL-based predicate filtering. We evaluate two pgvector-based baselines: Pg-IVF and Pg-HNSW, which use IVF-Flat and HNSW as the underlying vector indexes, respectively. Vectors and labels are stored in PostgreSQL tables with labels represented as INT[] arrays indexed using GIN for efficient containment queries. We use UNLOGGED tables to reduce index construction time. For Pg-IVF, we scan all vectors in the nprobenprobe buckets closest to the query vector and return the top-$k$ qualified vectors. Distances are computed for both qualified and unqualified vectors. Pg-HNSW employs iterative search that resumes exploration from the next $\mathit{ef}$ best candidates after each search iteration converges, terminating after finding $k$ qualified vectors or scanning at most $T_{max}$ vectors. To generate recall-throughput curves, we sweep nprobenprobe for Pg-IVF and $T_{max}$ for Pg-HNSW.
@@ 6.7. Ablation Study
Structural Constraints. Curator enforces structural sharing by constraining per-label indexes to follow the base index’s hierarchical partitioning, which potentially impacts performance since per-label indexes cannot freely adapt to label distributions. To evaluate this trade-off, we compare Curator against a variant where per-label indexes are constructed independently with the same hyper-parameters. Fig. 15(a) shows nearly identical performance between both versions, demonstrating that structural constraints have minimal impact on the quality of embedded per-label indexes. This finding is further supported by the similar performance between Curator with predicate indexed versus per-predicate indexing in § 6.2.
Vector-Label Correlation. Real-world datasets exhibit varying degrees of vector-label correlation—the tendency for vectors sharing the same label to cluster together in the embedding space due to semantic similarity. To assess Curator’s robustness across datasets with different correlation patterns, we test on a shuffled variant of YFCC-10M where labels are randomly reassigned, effectively eliminating vector-label correlation. Fig. 15(b) shows nearly identical recall-throughput trade-offs on both original and shuffled datasets, confirming Curator’s robust performance across varied data distributions. We additionally quantify correlation using the ”Query Correlation” metric from ACORN (Patel et al., 2024), which compares a query’s distance to the set of qualified vectors against a random set of equal size; larger values indicate stronger clustering by label. On YFCC-10M, the average query correlation is 0.161, whereas on the label-shuffled dataset it is 0.028, showing the effectiveness of label shuffling and confirming that the original dataset exhibits meaningful correlation.
Tree Traversal Overhead. Unlike graph-based indexes where all distance computations target individual vectors, Curator performs distance computations against both cluster centroids (during tree traversal) and vectors (during buffer scanning). To quantify the overhead of tree traversal, we measure the proportion of distance computations used for tree traversal versus buffer scanning. Fig. 15(c) shows that traversal overhead decreases from 26.9% to 18.9% as dataset size grows from 2M to 10M vectors, demonstrating that the overhead remains moderate and diminishes at larger scales.
@@ 8. Conclusion
Acknowledgements.
We thank the anonymous reviewers for their valuable feedback. This research was partially supported by NSF AI Institute on Edge Computing (Athena, grant No. 2112562) and NSF grants (2503010, 2402696, 2238665, 2402823), and by gifts from Adobe, Amazon, Meta, and IBM.